# Accelerating Residual Reinforcement Learning with Uncertainty Estimation

### v1

@@ Abstract
Residual Reinforcement Learning (RL) is a popular approach for adapting pretrained policies by learning a lightweight residual policy that provides corrective actions. While Residual RL is more sample-efficient than finetuning the entire base policy, existing methods struggle with sparse rewards and are designed for deterministic base policies. We propose two improvements to Residual RL that further enhance its sample efficiency and make it suitable for stochastic base policies. First, we leverage uncertainty estimates of the base policy to focus exploration on regions in which the base policy is not confident. Second, we propose a simple modification to off-policy residual learning that allows it to observe base actions and better handle stochastic base policies. We evaluate our method with both Gaussian-based and Diffusion-based stochastic base policies on tasks from Robosuite and D4RL, and compare against state-of-the-art finetuning methods, demo-augmented RL methods, and other residual RL methods. Our algorithm significantly outperforms existing baselines in a variety of simulation benchmark environments. We also deploy our learned polices in the real world to demonstrate their robustness with zero-shot sim-to-real transfer.
Figure 1: We propose two improvements to accelerate Residual RL : 1) We use an uncertainty metric to constrain exploration by only adding a residual for uncertain base actions. 2) We modify the off-policy critic to learn the $Q$ function for the combined action.
@@ I Introduction
Residual Reinforcement Learning is popular in robotics to improve the performance of pretrained policies by training a separate policy that outputs residual actions  [1, 2]. Directly fine-tuning the pretrained policy is often computationally expensive, especially in the case of policies with a large number of parameters [3, 4] and is prone to instability [5]. In contrast, Residual RL provides an efficient alternative to refine the base policy with minimal additional computation. This residual correction enables the agent to make targeted improvements, however the base policy was built.
@@
We evaluate our approach on a variety of manipulation tasks from Robosuite [8] and Franka Kitchen tasks from D4RL [11]. We test our approach on both Gaussian Mixture Model-based and Diffusion-based base policies. Finally, we compare our approach with a state-of-the-art finetuning method [12], a demo-augmented RL method [13], and other Residual RL methods [14, 1, 2]. Our proposed approach outperforms or is comparable to other baselines in all tasks. We also perform several ablation studies to test various aspects of our algorithm and also deploy the learned policies on a real robot to demonstrate sim-to-real transfer. Our proposed approach is visualized in Figure 1 with the main contributions summarized as follows:
@@ II Background
In Residual RL, we assume that we have a suboptimal base policy $\pi_{b}$. The objective is to learn a lightweight residual policy $\pi_{r}$, on top of the base policy that gives a corrective action $a_{r}$ to produce a more accurate and robust policy. Thus, the final policy executed in the environment is :
@@ III Related Work
Residual RL for Stochastic Base Policies. Residual RL, first introduced for robotics in [2] and [1], learns a corrective residual policy over a base controller, which can be either hand-designed or derived from model-predictive control. Importantly, these methods assume a deterministic base controller, as the residual policy is not conditioned on the base action. However, current state-of-the-art imitation learning algorithms like Diffusion policy [3] and GMM-based policies [8] are non-deterministic, making the original Residual RL formulation insufficient due to the lack of information about the base policy. Some works introduce noise in the base action to enhance robustness and induce stochasticity [6, 7] . Other works inform residual learning about the base policy by conditioning it on the learned bottleneck features of the base policy [9], and incorporating the base action in the observed state to inform the residual policy for on-policy learning [10]. In contrast, our work modifies off-policy RL to learn the $Q$ function for the combined action taken in the environment (i.e. the sum of base and residual action), and the actor uses the same $Q$ function to select a residual action. Therefore, our Residual RL formulation can handle the stochasticity of the base policies by making the base action observable to the critic while also being invariant to the split between the base action and residual action. Policy Decorator [14] also uses the combined action as an input to their critic for Residual RL, but we do a thorough quantitative evaluation with ablations to show that it significantly improves performance.
Imitation Learning and Residual RL. Several works have explored the integration of base policies trained using Imitation Learning (IL) with Residual RL. Residual RL has been applied to insertion tasks [15], where demonstrations are incorporated as an auxiliary behavior cloning (BC) loss during RL training [16]. FISH [17] employs a non-parametric base policy alongside a residual policy that uses optimal transport matching against offline demonstrations as the reward. BeT [18] introduces a residual action corrector that refines continuous actions on top of a discretized imitation policy. IBRL [13] does not directly use Residual RL, but it bootstraps an RL policy from an IL policy by using IL actions as alternative proposals for both online exploration and critic updates. Closest to our method is Policy Decorator [14], which learns bounded residual actions using controlled exploration. Unlike Policy Decorator, which uniformly samples action from the base policy and residual policy, we use uncertainty estimates of the base policy to decide when to learn corrective residual actions.
@@ IV-A Uncertainty Aware Residual RL
Prior works in Residual RL suffer from unrestrained exploration as they learn corrective residual actions uniformly over the entire state space. Our key insight is to improve exploration by focusing residual learning on regions in which the base policy is not confident. We propose using the uncertainty of the base policy to decide when to learn a residual action for the base policy. If the base policy is certain about its action for the current state, we directly use the action from the base policy $a_{b}$ in the environment; we only use a corrective residual action when the base policy is uncertain. Our proposed approach is agnostic to the uncertainty quantification method, and we demonstrate our method with two metrics, namely distance-to-data and ensemble variance. Distance-to-data has been used to calibrate the uncertainty of a model by measuring how out-of-distribution the current state is from an existing dataset [20]. For a dataset $D$ with each datapoint having $F$ features, we can estimate uncertainty using the minimum L2 distance of the current state $s$ to all points $d\in D$ :
$\text{uncertainty}_{d}(s)=\min_{d\in D}\sqrt{\sum_{i=1}^{F}(d_{i}-{\color[rgb]{0,0,0}s_{i}})^{2}}\ .$ (3)
Another popular approach to estimate uncertainty is by measuring the variance in predicted actions among an ensemble of policies. For an ensemble of $N$ base policies $\pi_{b}\in\pi_{B}$, the ensemble uncertainty can be defined as -
@@
Furthermore, we use an uncertainty threshold $\tau$ to measure the confidence of the base policy. This can be formulated as
@@ IV-B Optimizing Residual RL for Stochastic Policies
The original Residual RL algorithms are formulated to learn only using the partial residual actions, operating under the assumption that the underlying base policy is deterministic and can be implicitly inferred. Therefore, it learns the $Q$ function for only the partial residual action which is different from the action taken in the environment:
@@
Incorporating stochastic policies into the residual transition function $T_{r}$ can make it much harder to learn as they are noisier in their predictions and hence difficult to model. Thus, we suggest using Eq 2 to retreat back to the original MDP formulation. Consequently, the $Q$ function is learned for the combined action $a_{c}$, which is the actual action used during environment interaction :
@@
Previously, ResiP [10] proposed augmenting the observed state with base action to provide the missing information for on-policy RL. We instead propose learning the critic for the combined action $a_{c}$ for off-policy RL to provide the necessary information about the base policy to the $Q$ function while also making it invariant to the split between residual action and the base action. Specifically, we modify Soft Actor-Critic [28] in the following ways (changes marked in green). Initially, we store both the base action $a_{b}$ and the combined action $a_{c}$ in the replay buffer. While computing the target values, we should add the base action $a_{b}$ to the residual action $a_{r}$ sampled from the actor as follows:
$\begin{split}y{\color[rgb]{0,0,0}(r,s^{\prime},d)}&=r+\gamma(1-d)\ *\\ &\quad(\min_{i=1,2}Q_{\phi^{\prime}_{i}}(s^{\prime},{\color[rgb]{0,.5,.5}a^{\prime}_{r}+{a^{\prime}_{b}}})-\alpha\log\pi_{r}({\color[rgb]{0,.5,.5}a^{\prime}_{r}}|s^{\prime})),\\ &\quad{\color[rgb]{0,.5,.5}a^{\prime}_{r}\sim\pi_{r}(\cdot|s^{\prime})}\ .\end{split}$ (9)
While updating the $Q$ function, we should use the combined action stored in the replay buffer:
$J_{Q}(\phi_{i})=\mathbb{E}\left[\left(Q_{\phi_{i}}(s,{\color[rgb]{0,.5,.5}a_{c}})-y{\color[rgb]{0,0,0}(r,s^{\prime},d)}\right)^{2}\right],\quad i=1,2\ .$ (10)
When updating the actor, we can again add the base action to get the $Q$ value:
@@
The complete algorithm is described in Alg. 1 with the proposed changes in SAC marked in green.
@@ V Experiments
The setup for our experiments is described in Sec. V-A. We provide details on the baselines used in Sec. V-B and conduct an analysis of our results in Sec. V-C. We compare our approach with deterministic and stochastic base policy in Sec. V-D. We extend our approach to image-based environment in Sec V-E and perform real-world experiments in Sec V-F. Finally, we describe our ablation studies in Sec. V-G.
@@ V-A Experiment Setup
We define the robotic manipulation tasks for our experiments in Sec. V-A1. All environments use sparse rewards and state-based observations with task visualizations in Fig. 2. We evaluate all experiments with 5 seeds over 20 runs and provide hyperparameters in Appendix VI-C.
@@ V-A1 Environments
Robosuite: We evaluate the performance of the Panda robotic arm on three distinct tasks from the Robosuite simulator  [8]. We receive a sparse reward of 1 for successfully completing the task and 0 otherwise. We use the following tasks: 1) Lift task, where the robot arm must pick up a block placed at a random initial position on the table. 2) Can task, where the robot arm has to pick up a can from one table and place it in the top right corner of another table. 3) Square task, where the robot arm must pick up a square nut from the table and place it onto a square bolt.
@@ V-A2 Base Policies
We consider two kinds of IL base policies to test the robustness of our algorithm :
@@
Diffusion-based policy: To ensure a consistent comparison, we use the same Diffusion policies from DPPO [12] for our experiments. For Robosuite tasks, they use noisier multi-human demonstrations, while for the Franka Kitchen environments, they directly use datasets from D4RL benchmark.
@@ V-B Baselines
We compare our proposed approach against finetuning methods, demo augmented RL methods, and other Residual RL methods.
Finetuning methods: As a baseline for Diffusion-based policies, we use the recently proposed DPPO [12]. DPPO formulates the denoising process of the Diffusion policy as a separate MDP, effectively modeling the entire trajectory as a sequence of MDPs. The policy is then optimized using policy gradient across this entire chain of MDPs.
@@
Residual RL methods: We test our method against the most closely related approach and current state-of-the-art residual learning algorithm, Policy Decorator [14], which also aims to mitigate excessive exploration in Residual RL. It addresses this by sampling actions uniformly from both the residual and base policies while gradually decreasing the proportion of actions taken from the base policy. Additionally, it bounds the actions of the residual policy. For completeness, we also compare our approach with the standard Residual RL [1, 2] algorithm, incorporating our proposed modifications to the critic.
@@ V-C Results and Analysis
We analyze the results for GMM-based policies in Sec. V-C1, and the results of Diffusion policies in Sec. V-C2. We compared the performance of our method with both distance-to-data and ensemble variance as the uncertainty metric for these results.
@@ V-C1 GMM-based policies
We present the results of our experiments with GMM-based policies in Figure 3. We plot the success rate over the course of environment interactions. Our method with distance to data metric (red line) outperforms all the baselines in the three Robosuite environments. We note that there is an initial dip in the performance for our method where the residual policy is in the exploration phase, but it starts improving once the exploration phase ends. IBRL performs the best out of the other baselines with its performance comparable to our method for the Lift Task, though the initial dip in performance is more significant, and it is still unstable afterwards. We performed a hyperparameter sweep for the two additional parameters of Policy Decorator, namely the residual bound and the decay rate. Our method converges in the same number of timesteps while Policy Decorator’s performance is still improving. We suspect it is due to the more targeted exploration of our method using uncertainty estimates. The standard Residual Policy Learning method only goes over the base policy performance for the Lift Task.
@@ V-C2 Diffusion policies
We present the results for Diffusion policies in Figure 4. We run the experiments for the Kitchen Complete environment of the D4RL benchmark. In Robosuite, we perform experiments in the Can and Square environments, excluding the Lift environment because of the near-optimal performance of Diffusion policy on that task. We performed a hyperparameter sweep of Policy Decorator as mentioned in Sec. V-C1. Similarly, our method is able to outperform Policy Decorator in all environments. Our approach is able to achieve higher success rates than all the baselines in the Kitchen Complete environment. We note that even though DPPO’s performance is stable in Kitchen Complete and Can environments, its improvement over the initial performance of the base policy is slow compared to our approach, despite an initial dip in the performance. Our method with ensemble variance as a metric has comparable performance to DPPO for the Square task. These results suggest that our approach is most promising in scenarios where the initial base policy performance is average, and is comparable in scenarios where the initial base policy performance is bad. Also, ensemble variance as a metric works better or is comparable to distance-to-data for the Can and Square environment but not in the Kitchen Complete environment. We hypothesize that this is because of the good quality of demonstrations in the Kitchen Complete environment while the noisier multi-human demonstrations of Can and Square tasks can result in a noisier distance-to-data metric, though our method is still able to converge for the same.
Figure 5: Learning with combined action and residual action for stochastic and deterministic base policies.
Figure 6: Results on Can-Image Task.
V-D Combined Action vs Residual Action
To emphasize the significance of utilizing combined actions for stochastic policies, we compared our modified SAC algorithm for Residual RL without uncertainty estimates to the original Residual RL formulation. For this comparison, we used the GMM policy as the stochastic base policy and a standard MLP policy as the deterministic base policy, applied to the Lift task in Robosuite. The results of using combined actions and residual actions during learning for both stochastic and deterministic base policies are shown in Figure 5. The findings reveal that relying solely on the residual action does not yield effective results for stochastic base policies, highlighting the necessity of combining actions in such cases. In contrast, for deterministic base policies, either residual actions or combined actions can be effectively used.

### v2

@@ Abstract
Residual Reinforcement Learning (RL) is a popular approach for adapting pretrained policies by learning a lightweight residual policy that provides corrective actions. While Residual RL is more sample-efficient than finetuning the entire base policy, existing methods struggle with sparse rewards and are designed for deterministic base policies. We propose two improvements to Residual RL that further enhance its sample efficiency and make it suitable for stochastic base policies. First, we leverage uncertainty estimates of the base policy to focus exploration on regions in which the base policy is not confident. Second, we propose a simple modification to off-policy residual learning that allows it to observe base actions and better handle stochastic base policies. We evaluate our method with both Gaussian-based and Diffusion-based stochastic base policies on tasks from Robosuite and D4RL, and compare against state-of-the-art finetuning methods, demo-augmented RL methods, and other Residual RL methods. Our algorithm significantly outperforms existing baselines in a variety of simulation benchmark environments. We also deploy our learned policies in the real world to demonstrate their robustness with zero-shot sim-to-real transfer. Paper homepage : lakshitadodeja.github.io/uncertainty-aware-residual-rl/
publicationid: pubid: 0000–0000/00\$00.00 © 2021 IEEE
Figure 1: We propose two improvements to accelerate Residual RL : 1) We use uncertainty estimation to constrain exploration around the base policy. 2) We modify the off-policy critic to learn the $Q$ function for the combined action. We test our method with two uncertainty metrics, i.e. distance to data and ensemble variance.
@@ I Introduction
Residual Reinforcement Learning improves the performance of pretrained policies by training a separate policy to output corrective actions  [26, 14]. Directly fine-tuning the pretrained policy is often computationally expensive, especially in the case of policies with a large number of parameters [7, 4], and is prone to instability [18]. In contrast, Residual RL provides an efficient alternative to refine the base policy with minimal additional computation. This residual correction enables the agent to make targeted improvements however the base policy was built.
@@
We evaluate our approach on a variety of manipulation tasks from Robosuite [17] and Franka Kitchen tasks from D4RL [8]. We test our approach on both Gaussian Mixture Model and Diffusion base policies. Finally, we compare our approach with a state-of-the-art finetuning method [21], a demo-augmented RL method [13], and other Residual RL methods [29, 26, 14]. Our proposed approach outperforms or is comparable to other baselines in all tasks. We also perform several ablation studies to test various aspects of our algorithm and deploy the learned policies on a real robot to demonstrate sim-to-real transfer. Our proposed approach is visualized in Figure 1 with the main contributions summarized as follows:
@@ II Background
In Residual RL, we assume that we have a suboptimal base policy $\pi_{b}$. The objective is to learn a lightweight residual policy $\pi_{r}$, on top of the base policy that gives a corrective action $a_{r}$, producing a more accurate and robust combined policy. Thus, the final policy executed in the environment is:
@@ III Related Work
Residual RL for Stochastic Base Policies. Residual RL, first introduced for robotics in [14] and [26], learns a corrective residual policy over a base controller, which can be either hand-designed or derived from model-predictive control. Importantly, these methods assume a deterministic base controller, as the residual policy is not conditioned on the base action. However, current state-of-the-art imitation learning algorithms like Diffusion policy [7] and GMM-based policies [17] are non-deterministic, making the original Residual RL formulation insufficient due to the lack of information about the base policy. Some works introduce noise in the base action to enhance robustness and induce stochasticity [16, 30]. Other works inform residual learning about the base policy by conditioning it on the learned bottleneck features of the base policy [1], and incorporating the base action in the observed state to inform the residual policy for on-policy learning [3]. In contrast, our work modifies off-policy RL to learn the $Q$ function for the combined action taken in the environment (i.e. the sum of base and residual action), and the actor uses the same $Q$ function to select a residual action. Therefore, our Residual RL formulation can handle the stochasticity of the base policies by making the base action observable to the critic while also being invariant to the split between the base action and residual action. Policy Decorator [29] also uses the combined action as an input to their critic for Residual RL, but we do a thorough quantitative evaluation to show that it significantly improves performance.
Imitation Learning and Residual RL. Several works have explored the integration of Imitation Learning (IL) base policies with Residual RL. Residual RL has been applied to insertion tasks [23], where demonstrations are incorporated as an auxiliary behavior cloning (BC) loss during RL training [19]. FISH [12] employs a non-parametric base policy alongside a residual policy that uses optimal transport matching against offline demonstrations as the reward. BeT [25] introduces a residual action corrector that refines continuous actions on top of a discretized imitation policy. IBRL [13] does not directly use Residual RL, but it bootstraps an RL policy from an IL policy by using IL actions as alternative proposals for both online exploration and critic updates. Closest to our method is Policy Decorator [29], which learns bounded residual actions using controlled exploration. Unlike Policy Decorator, which uniformly samples actions from the base and residual policies, we use uncertainty estimates of the base policy to decide when to learn and apply corrective residual actions.
@@ IV-A Uncertainty Aware Residual RL
[…]
$\text{uncertainty}_{d}(s)=\min_{d\in D}\sqrt{\sum_{i=1}^{F}(d_{i}-s_{i}})^{2}\ .$ (3)
Another popular approach to estimating uncertainty is measuring the variance in predicted actions among an ensemble of policies. For an ensemble of $N$ base policies $\pi_{b}\in\pi_{B}$, the ensemble uncertainty can be defined as:
@@
We can use our desired uncertainty metric with a threshold $\tau$ to measure the confidence of the base policy. This can be formulated as
@@ IV-B Optimizing Residual RL for Stochastic Policies
The original Residual RL algorithms are formulated to learn only using partial residual actions, operating under the assumption that the underlying base policy is deterministic and can be implicitly inferred. Therefore, it learns the $Q$ function for only the partial residual action, which is different from the action taken in the environment:
@@
Incorporating stochastic policies into the residual transition function $T_{r}$ can make it much harder to learn, as they are noisier in their predictions and hence difficult to model. Thus, we suggest using Eq. 2 to retreat back to the original MDP formulation. Consequently, the $Q$ function is learned for the combined action $a_{c}$, which is the actual action used during environment interaction :
@@
Previously, ResiP [3] proposed augmenting the observed state with base action to provide the missing information for on-policy RL. We instead propose learning the critic for the combined action $a_{c}$ for off-policy RL, providing the necessary information about the base policy to the $Q$ function while also making it invariant to the split between the residual and base actions. Specifically, we modify Soft Actor-Critic [11] in the following ways (changes marked in green). Initially, we store both the base action $a_{b}$ and the combined action $a_{c}$ in the replay buffer. While computing the target values, we add the base action $a_{b}$ to the residual action $a_{r}$ sampled from the actor:
$\begin{split}y(r,s^{\prime},d)&=r+\gamma(1-d)\ *\\ &\Big[\min_{i=1,2}Q_{\phi^{\prime}_{i}}(s^{\prime},{\color[rgb]{0,.5,.5}a^{\prime}_{r}+{a^{\prime}_{b}}})-\alpha\log\pi_{r}({\color[rgb]{0,.5,.5}a^{\prime}_{r}}|s^{\prime})\Big],\\ &\quad{\color[rgb]{0,.5,.5}a^{\prime}_{r}\sim\pi_{r}(\cdot|s^{\prime})}.\end{split}$ (9)
When updating the $Q$ function, we use the combined action stored in the replay buffer:
$J_{Q}(\phi_{i})=\mathbb{E}\left[\left(Q_{\phi_{i}}(s,{\color[rgb]{0,.5,.5}a_{c}})-y(r,s^{\prime},d)\right)^{2}\right],\quad i=1,2\ .$ (10)
When updating the actor, we again add the base action to get the $Q$ value:
@@
The complete algorithm is described in Alg. 1 with the proposed changes from SAC marked in green.
@@ V Experiments
In this section we describe our experiment setup, baselines, results and ablations.
@@ V-A1 Environments
We conduct evaluations on the environments visualized in Fig. 2 and described below. All environments use sparse rewards and state-based observations.
Robosuite: We evaluate on three tasks from the Robosuite simulator [17]: 1) Lift, where the robot arm must pick up a block placed at a random initial position on the table. 2) Can, where the robot arm has to pick up a can from one table and place it in the top right corner of another table. 3) Square, where the robot arm must pick up a square nut from the table and place it onto a square bolt.
@@ V-A2 Base Policies
We consider two kinds of IL base policies to test the robustness of our algorithm.
@@
Diffusion-based policy: We also test our method with a diffusion base policy [7]. To ensure a consistent comparison, we use the same Diffusion policies from DPPO [21] for our experiments. For Robosuite tasks, they use noisier multi-human demonstrations, while for the Franka Kitchen environments, they directly use datasets from D4RL.
@@ V-B Baselines
Finetuning methods: As a baseline for Diffusion-based policies, we use the recently proposed DPPO [21]. DPPO formulates the denoising process of the Diffusion policy as a separate MDP, effectively modeling the entire trajectory as a sequence of MDPs. The policy is then optimized using policy gradients across this entire chain of MDPs.
@@
Residual RL methods: We test our method against the most closely related approach and current state-of-the-art residual learning algorithm, Policy Decorator [29], which also aims to mitigate excessive exploration in Residual RL. It addresses this by sampling actions uniformly from both the residual and base policies with a progressive exploration schedule. Additionally, it bounds the actions of the residual policy. For completeness, we also compare our approach with the standard Residual RL [26, 14] algorithm, incorporating our proposed modifications to the critic without using uncertainty estimation.
@@ V-C Results and Analysis
We evaluate all experiments with 5 seeds over 20 runs and provide hyperparameters in Appendix C. Hyperparameters.
@@ V-C1 GMM-based policies
We present the results of our experiments with GMM-based policies in Figure 3. We plot the success rate over the course of environment interactions. Our method with distance to data metric (red line) outperforms all the baselines in the three Robosuite environments. We note that there is an initial dip in the performance for our method where the residual policy is in the exploration phase, but it starts improving once the exploration phase ends. IBRL performs the best out of the other baselines, reaching comparable performance to our method for the Lift Task, though the initial dip in performance is more significant, and it is still unstable afterwards. We performed a hyperparameter sweep for the two additional parameters of Policy Decorator, namely the residual bound and the decay rate with more details in the Appendix D. Our method converges in the same number of timesteps, while Policy Decorator’s performance is still improving. We suspect it is due to the more targeted exploration of our method using uncertainty estimates. The standard Residual Policy Learning method only improves over the base policy performance for the Lift Task.
@@ V-C2 Diffusion policies
We present the results for Diffusion policies in Figure 4. We run the experiments for the Kitchen Complete environment of the D4RL benchmark. In Robosuite, we perform experiments in the Can and Square environments, excluding the Lift environment because of the near-optimal performance of Diffusion policy on that task. We performed a similar hyperparameter sweep of Policy Decorator as mentioned in Sec. V-C1 with details in Appendix D. Similarly, our method is able to outperform Policy Decorator in all environments. Our approach is able to achieve higher success rates than all the baselines in the Kitchen Complete environment with both types of uncertainty metrics. We note that even though DPPO’s performance is stable in Kitchen Complete and Can environments, its improvement over the initial performance of the base policy is slow compared to our approach, despite an initial dip in the performance. Our method with the ensemble variance metric has comparable performance to DPPO for the Square task. These results suggest that our approach is most promising when the initial base policy performance is average, and is comparable in scenarios where the initial base policy performance is bad. Also, we note that distance-to-data as a metric works better than ensemble variance for the Kitchen Complete environment but not the Can and Square environment. We hypothesize that this is because of the high quality demonstrations in the Kitchen Complete environment, whereas the noisier multi-human demonstrations of Can and Square tasks result in a noisier distance-to-data metric. Nevertheless, our method converges for both the metrics.
V-D Combined Action vs. Residual Action
To emphasize the significance of utilizing combined actions for stochastic policies, we compared our modified SAC algorithm for Residual RL without uncertainty estimates against the original Residual RL formulation. For this comparison, we used the GMM policy as the stochastic base policy and a standard MLP policy as the deterministic base policy, applied to the Lift task in Robosuite. The results for both stochastic and deterministic base policies are shown in Figure 5. The findings reveal that relying solely on the residual action does not yield effective results for stochastic base policies, highlighting the necessity of using our combined action formulation in such cases. In contrast, for deterministic base policies, either residual actions or combined actions can be used.
Figure 5: Learning with either the residual or the complete action works well with deterministic base policies, but learning with complete action is required for stochastic base policies.
Figure 6: Results on the Can-Image Task.